Solve the system using elimination. Write the solution as an ordered pair.
x + 2y = -7
3x - 4y = -1

Answer by stanbon(75887)

Solve the system using elimination. Write the solution as an ordered pair.
x + 2y = -7
3x - 4y = -1
------------
Modify::
2x + 4y = -14
3x - 4y = -1
-------
Add and solve for "x"::
5x = -15
x = -3
------
Solve for "y"::
x + 2y = -7
-3 + 2y = -7
2y = -4
y = -2
---------------
Cheers,
Stan H.
--------
--------------